OpenStudy (harley1312):

Will Give Medal 1. Kevin ran 4 miles more than Steve ran. The sum of their distances is 26 miles. How far did Steve run? The domain of the solution is . (a) Define your variable. (b) Write algebraic expressions for the distances each person ran. (c) Write an equation that can help you solve the problem. (d) Solve the equation. (e) Check your answer.

OpenStudy (harley1312):

domain of the solution is .

OpenStudy (princeharryyy):

let kevin is K and steven is S, ATQ, S = K +4 again S + K =26 solve both equations find S and K.

OpenStudy (harley1312):

{0,4,11,22} is the domain

OpenStudy (princeharryyy):

hold on.... it's K = 4 + S not S = K + 4

OpenStudy (princeharryyy):

How did u got that domain @Harley1312

OpenStudy (princeharryyy):

Ok. hold on. let me check.

OpenStudy (harley1312):

Okay

OpenStudy (princeharryyy):

your variables are S and K, Ok and they are in kms as it's the distance. Is that fine for the first one.

OpenStudy (harley1312):

What?? :|

OpenStudy (princeharryyy):

Now, for the second one. you see Kevin always runs 4 kms more than Steve, Right. So, Algebraic equation for Kevin is >> K = 4 +s on the other hand steve can't run more than 22 kms as the sum of both of them covering the distance is 26kms, as is given in the question. So the equation for Steve becomes>> S = 26-K, with the restriction that S can not be greater than 22 as K = S +4. Does that make sense to you. If it does we can proceed further.

OpenStudy (princeharryyy):

Sorry there are in miles not kms, didn't see that.

OpenStudy (princeharryyy):

they*

OpenStudy (princeharryyy):

does that make sense to you @Harley1312

OpenStudy (harley1312):

Yes, sorry I was filling in my paper.

OpenStudy (princeharryyy):

just go through it and let me know.

OpenStudy (harley1312):

I got it down

OpenStudy (princeharryyy):

Ok, so Shall we proceed with part 3

OpenStudy (harley1312):

Yes :)

OpenStudy (princeharryyy):

it asked for equation in third part we already did it. let's check the fourth part.

OpenStudy (harley1312):

Yeah I already filled in the third part , sorry i forgot to tell you

OpenStudy (princeharryyy):

ok. Np.

OpenStudy (harley1312):

Okay

OpenStudy (princeharryyy):

hold on doing the 4th part on paper will let you know :)

OpenStudy (harley1312):

Okay hun thank you

OpenStudy (princeharryyy):

hey, would you mind telling me something.

OpenStudy (harley1312):

No ?

OpenStudy (harley1312):

If it's a personal question message me

OpenStudy (princeharryyy):

NO no, It's about the question itself.

OpenStudy (harley1312):

Okay, ask away..

OpenStudy (princeharryyy):

Is that domain given in the question itself or you wrote that.

OpenStudy (harley1312):

It's in the question

OpenStudy (princeharryyy):

Just try and understand what I am trying to make you understand, Ok. Can you do that for me?

OpenStudy (princeharryyy):

@harley1312

OpenStudy (princeharryyy):

As you know sum of the distance of kevin and steve is 26 miles and kevin walks 4 miles more than sam, i.e k = s +4 and k+s = 26 miles. Now if, k walks 26 miles that means steve must have walked 22 kms as k = s +4 but than sum of k and s becomes 26 +22 = 48 miles that means kevin didn't walk 26 kms...

OpenStudy (princeharryyy):

Again if steve walks 0 km than kevin will walk only 4 kms as k = s +4 but than sum of kevin and steve would be 0 +4 = 4kms which is not equal to 26 kms

OpenStudy (princeharryyy):

Now if steve walks 11 miles than kevin must have walked 15 miles as k = s +4 => k = 11+4 = 15 miles also k+s in this case becomes 15 +11 = 26 miles, hence steve walks 11 miles and kevin 15 miles

OpenStudy (princeharryyy):

steve can't walk 22 miles as k = s+4 becomes 22+4 = 26miles but than k+s can not be 26 miles as k alone has walked 26miles and s 22 miles adding up to 48 miles. So, Steve must have walked 11 miles and kevin 15 miles. I hope you understand that. for the solving of equations => K = S+4, say it equation 1 K+S =26, say it equation 2 Substitute K in equation 2 becomes, S +4 +S = 26, 2S = 22 and S= 11 miles answer. @harley1312